solve 4x+9y+8z=14 x+2y+5z=0 x-4y-3z=-4
Found 2 solutions by amfagge92, MathTherapy:
Answer by amfagge92(93)

4X+9Y+8Z=14---i
X+2Y+5Z=0.....ii
X-4Y-3Z=-4....iii
SOLVING
i-ii:y-2z=14...iv
iii-ii:-6y=-4+8z..v
from eqn iv
y=14+2z----vi
put y in eqn v
-6(14+2z)=-4+8z
-84-12z=-4+8z
-80=20z
z=-4
from eqn vi
y=14+2z
y=14+2(-4)
y=6
X Can be found using any of the initial eqn
x-4y-3z=-4
x-4(6)-3(-4)=-4
x-24+12=-4
X=8.
therefore,X=8 ,Y=6 ,Z=-4.

Answer by MathTherapy(10552)


solve 4x+9y+8z=14 x+2y+5z=0 x-4y-3z=-4
Correct answer: 
Ignore all other INCORRECT answers.